OpenStudy (anonymous):

Solve 2x2 + 3x + 6 = 0. Round solutions to the nearest hundredth. x ≈ −2.64 and x ≈ 1.14 x ≈ −1.14 and x ≈ 2.64 No real solutions x ≈ −1.11 and x ≈ −4.89

OpenStudy (anonymous):

I think its No real solutions... Am i right?

OpenStudy (cwrw238):

the discriminant b^2 - 4ac = 9 - 4*2*6 = -39 - right

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (cwrw238):

yw